OpenStudy (anonymous):

The force required to stretch a spring beyond its natural length is proportional to how far the spring is being stretch.If a force of 8 pounds stretchs a spring 16 centimeters beyond its natural length,what force,in pounds,is need to stretch this spring 42 centimeter beyond its natural length. RATIO PROBLEM A) 15 B)21 C)31 D)37 E)42

OpenStudy (anonymous):

f = k e where f = force, k = constant e = extension( difference in length from initial) 8 = k *16 k = 1/2 f = k * 42 f = 1/2 * 42 f = 21pound

OpenStudy (anonymous):

explain please

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

a spring always follows a constant extension that, is for each amount of force supplied, the spring extends by a constant amount, say when u give a force of 1 N it extends by 2cm, then if u give a force of 2N it extends by 4 cm and so on, so we can set up a linear equation where force and extension can be used to find the constant Force = Constant * Extension (now we know the force as well as the extension, so we can substitute and find the constant) then we can use that constant to find the force for a given extension) lets see how it works 8 pounds = constant * 16 cm(extension) constant = 1/2 ( this is the way it extends ) now use this constant to find force for 42 cm extension F = 1/2 * 42 F = 21 pounds